Spin Spinner 3, then Spinner 2. What is the probability of getting Foward, followed by Yellow or Green?
Spinner 2 has (3 equal areas) (Green,Yellow,Red)
Spinner 3 has (2 equal areas) (Foward and Backward)

Answer by Boreal(15235)

Independent, so multiply probabilities
P(F)=1/2
P(Y or G)=2/3
(1/2)(2/3)=1/3.